When the car park at Brinsons empties at the end of the day, the probability that a car turns left at the first junction is 3/4 and that a car turns left at its NEXT junction is 2/5. A car leaves the car park. What is the probability that
a) It turns right and then right again
b) It makes one left turn and one right turn at the first two junctions it encounters.

Answer by stanbon(75887)

When the car park at Brinsons empties at the end of the day, the probability that a car turns left at the first junction is 3/4 and that a car turns left at its NEXT junction is 2/5. A car leaves the car park. What is the probability that
a) It turns right and then right again: (1/4)(3/5) = 3/20
-----------------------
b) It makes one left turn and one right turn at the first two junctions it encounters.
(3/4)(3/5) + (1/4)(3/5) = 9/20 + 3/20 = 12/20 = 3/5
==================
Cheers,
Stan H.
==================